OpenStudy (anonymous):

using the conversion box below, find how may quarts are in 384 ounces. Measurement : 1 quart = 4 cups and 1 cup = 8 ounces

OpenStudy (kayders1997):

1 quart=4cups?

OpenStudy (kayders1997):

For this one change oz to cups so 384 oz. 8 oz ----------- = cross multiply for x X. 1 cup

OpenStudy (kayders1997):

Sorry idk how to make it look neat

OpenStudy (kayders1997):

So basically your doing 384=8x so to solve for x you have to? And this is for cups

OpenStudy (kayders1997):

Do you have options? Just to see if one of the options is what I got

OpenStudy (anonymous):

a= 1/32 b=1/12 c=12 d=32

OpenStudy (anonymous):

by the way thank you for the help

OpenStudy (kayders1997):

No problem once you get that which should be 48 divide that by 4 because you have to get to quarts and you should get 12

OpenStudy (kayders1997):

Hopefully this makes some sense? Of going from oz to cups and cups and quarts

OpenStudy (anonymous):

yes i did make sense..thanks